# Interpretability-Aware Pruning for Efficient Medical Image Analysis

### v1

@@ Abstract
Keywords:
Model Pruning, Model Compression, Explainable AI (XAI), Trustworthy AI in Healthcare, AI Alignment.
@@ 4.1 Datasets and Models
We utilized four diverse medical imaging datasets, each with unique clinical contexts:
@@
We applied our pruning methodology across three widely used architectures: VGG19 [28], ResNet50 [29], and Vision Transformer (ViT) [30]. Each model was trained in a supervised fashion on all datasets to establish a strong baseline (see Section 0.B). This process incorporated data augmentation techniques, learning rate schedulers, optimization algorithms, and loss functions, including Cross Entropy, to achieve optimal results.
@@ 4.2 Pruning Process and Evaluation
For each model and dataset, we followed the methodology outlined in Section 3. Importance scores were aggregated from a subset of samples obtained from the specified sampling techniques (most Confidence, Random, Clustering with 10 samples per class from the training set). Neurons with the lowest aggregated importance scores were pruned based on a defined pruning rate. We evaluated the models’ accuracy and computational efficiency at various pruning rates, specifically observing the point of "lossless pruning" where accuracy is maintained despite significant compression.
@@ 5 Results and Discussion
The results demonstrate the effectiveness of the proposed pruning methodology across different medical imaging datasets and model architectures. This section highlights core findings related to model resilience under high pruning rates, the impact of attribution methods, and the influence of sampling strategies.
Table 1 and Figure 2 summarize pruning performance across the experimental spectrum. Across all datasets and models, interpretability-guided pruning using LRP, DLB, and IG consistently enables high model compression while maintaining accuracy within a 5% margin, even at moderate pruning rates.
@@ 5.1 Model-Wise Observations
We observed across all datasets that, in most cases, ViTs seem to have the highest accuracy followed by Vgg and Resnet. Among the architectures, ViTs demonstrate exceptional robustness, consistently preserving performance even beyond 65–80% pruning. Notably, in certain configurations (e.g., CPN with LRP or DLB), ViTs withstand pruning rates of 85% with minimal degradation. This performance suggests a higher degree of redundancy or feature richness in transformer-based models compared to CNNs. VGG architectures typically tolerate up to 45–55% pruning with less than 5% accuracy drop, whereas ResNets are comparatively sensitive, maintaining stability only up to 25–35% pruning in most settings. The observed variation highlights the architectural differences: residual connections in ResNet may cause greater entanglement of information across neurons, making unstructured neuron pruning comparatively less impactful.
@@ 5.3 Impact of different sampling strategies
Clustering-based sampling performs best on average across attribution methods, particularly noticeable at moderate pruning levels (30–50%). As shown in Table 1, LRP with Clustering shows the lowest drop, outperforming Random and Confidence-based alternatives. Confidence-based sampling performs competitively, particularly for IG and DLB, where it occasionally surpasses Random selection in maintaining model stability.
Interestingly, in certain configurations (especially at lower pruning rates like 15–30%), a slight increase in accuracy is observed. This counterintuitive behavior arises because pruning also removes neurons that may have a negative or noisy influence on predictions. By discarding these misleading neurons, the model generalizes better, resulting in a net accuracy gain.
The observed high variance is expected as these results are averaged across different datasets and architectures; however, overall trends remain consistent.
@@ 6 Implications for Medical Image Analysis
7 Conclusion

### v2

@@ 4.1 Datasets and Models
We utilized four medical imaging datasets, each with unique clinical contexts:
@@
We applied our pruning methodology across three widely used architectures: VGG19 [27], ResNet50 [14], and Vision Transformer (ViT) [5]. Each model was trained in a supervised fashion on all datasets to establish a strong baseline. This process incorporated data augmentation techniques, learning rate schedulers, optimization algorithms, and loss functions, including Cross Entropy, to achieve optimal results.
@@ 4.2 Pruning Process and Evaluation
For each model and dataset, we followed the methodology outlined in Section 3. Importance scores were aggregated from a subset of samples obtained from the specified sampling techniques (most Confidence, Random, Clustering with 10 samples per class from the training set). Neurons with the lowest aggregated importance scores were pruned based on a defined pruning rate. We evaluated the models’ accuracy and computational efficiency at various pruning rates, specifically observing the point of "lossless pruning" where accuracy is maintained despite significant compression. Although the importance score computation incurs a one-time overhead proportional to model size, this cost is theoretically justified by the resulting gains in post-pruning inference efficiency, thereby amplifying benefits in resource-limited clinical settings.
@@ 5 Results and Discussion
The results demonstrate the effectiveness of the proposed pruning methodology across different medical imaging datasets and model architectures. This section highlights core findings related to model resilience under high pruning rates, the impact of attribution methods, and the influence of sampling strategies. Table 1 and Figure 2 summarize pruning performance across the experimental spectrum. Across all datasets and models, interpretability-guided pruning using LRP, DLB, and IG consistently enables high model compression while maintaining accuracy within a 5% margin, even at moderate pruning rates.
@@ 5.1 Model-Wise Observations
We observed across all datasets that, in most cases, ViTs seem to have the highest accuracy followed by Vgg and Resnet. Among the architectures, ViTs demonstrate exceptional robustness, consistently preserving performance even beyond 65–80% pruning. Notably, in certain configurations (e.g., CPN with LRP or DLB), ViTs withstand pruning rates of 85% with minimal degradation. This performance suggests a higher degree of redundancy or feature richness in transformer-based models compared to CNNs. VGG architectures typically tolerate up to 50% pruning with less than 5% accuracy drop, whereas ResNets are comparatively sensitive, maintaining stability only up to 30% pruning in most settings. The observed variation highlights the architectural differences: residual connections in ResNet may cause greater entanglement of information across neurons, making unstructured neuron pruning comparatively less impactful.
@@ 5.3 Impact of different sampling strategies
Clustering-based sampling performs best on average across attribution methods, particularly noticeable at moderate pruning levels (30–50%). As shown in Table 1, LRP with Clustering shows the lowest drop, outperforming Random and Confidence-based alternatives. Confidence-based sampling performs competitively, particularly for IG and DLB, where it occasionally surpasses Random selection in maintaining model stability. Interestingly, in certain configurations (especially at lower pruning rates like 15–30%), a slight increase in accuracy is observed. This counterintuitive behavior arises because pruning also removes neurons that may have a negative or noisy influence on predictions. By discarding these misleading neurons, the model generalizes better, resulting in a net accuracy gain. The sampling strategy also influences pruning stability; while minor fine-tuning may be required for specific datasets, the overall framework remains robust across sampling choices. The observed high variance is expected as these results are averaged across different datasets and architectures; however, overall trends remain consistent.
6 Conclusion
@@ 7 Implications for Medical Image Analysis
Although further empirical validation is needed, we hypothesize that pruning pre-trained models using region-specific, representative data could improve generalization at deployment. This approach may mitigate domain shift and preserve model reliability in real-world clinical settings.
Disclosure of Interests
The authors are affiliated with AryaXAI.com, part of Aurionpro Solutions Limited. DL-Backtrace was introduced by AryaXAI. The authors declare no further competing interests.